OpenStudy (carolinar7):

The temperature of a chemical reaction ranges between −10 degrees Celsius and 50 degrees Celsius. The temperature is at its lowest point when t = 0, and the reaction completes 1 cycle during a 6-hour period. What is a cosine function that models this reaction?

OpenStudy (carolinar7):

f(t) = −30 cos pi over 3t + 20 f(t) = −60 cos pi over 3t + 30 f(t) = 30 cos 6t + 20 f(t) = 60 cos 6t + 30

OpenStudy (carolinar7):

@laughoutloud

OpenStudy (carolinar7):

Module 10 exam

OpenStudy (carolinar7):

@mathway

OpenStudy (carolinar7):

@Michele_Laino

OpenStudy (laughoutloud):

idk the answer I finished everything last night and can't get back in

OpenStudy (laughoutloud):

sorry

OpenStudy (carolinar7):

Oh can you explain it though

OpenStudy (carolinar7):

@Nnesha

OpenStudy (laughoutloud):

@phi

OpenStudy (carolinar7):

@dan815

OpenStudy (carolinar7):

@pooja195

OpenStudy (carolinar7):

I guessed it and got it right